The following table shows the Myers-Briggs personality preferences for a random sample of 406 people in the listed professions. E refers to extroverted and I refers to introverted. Row Total 107 162 137 406 Occupation all denominations 47 90 84 185221 60 72 53 M.D. Lawyer Use the chi-square test to determine if the listed occupations and personality preferences are independent at the 0.05 level of significance. (a) What is the level of significance? State the null and alternate hypotheses. O Ho: Myers-Briggs preference and profession are not independent H: Myers-Briggs preference and profession are not independent. O Ho: Myers-Briggs preference and profession are independent H: Myers-Briggs preference and profession are independent. Ho: Myers-Briggs preference and profession are independent H: Myers-Briggs preference and profession are not independent. Ho: Myers-Briggs preference and profession are not independent H: Myers-Briggs preference and profession are independent. Find the value of the chi-square statistic for the sample. (Round the expected frequencies to at least three decimal places. Round the test statistic to three decimal places.) Are all the expected frequencies greater than 5? O Yes O No
Answer #1

b)

Obs fre (f) Exp freq e) Total Occupation E Clergy MD Lawyer Total Total 60 72 53 185 47 90 84 221 107 162 137 406 48.756 73.8

The test statistic is

\chi ^{2} = \sum \frac{(f-e)^2}{e} =7.461